In the right triangle find the length of the side not given Give an exact answer and an approxiimation to three decimal places..
Side A=3 B=5 and C=
Pleas show each step

Answer by edjones(8007)

a^2+b^2=c^2
3^2+5^2
=9+25
=34
sqrt(34)=5.831
.
Ed